1.1 Suppose you buy a round lot of MM Industries stock on 55 percent margin when the stock is selling at $20 a share. The broker charges a 10 percent annual interest rate, and commissions are 3 percent of the total stock value on both the purchase and sale. A year later, you receive a $0.50 per share dividend and sell the stock for 27. What is your rate of return on the investment?


1
Expert's answer
2022-03-30T12:56:01-0400

Cost of share =$20

Margin = 55%\times 20 =11

Rest is the loan from broker= 20-11=9

Total repayment after 1 year @10% interest rate= 9\times(1+0.1)=9.9